OpenStudy (anonymous):

Can someone help me find cos 2theta if sin (theta)=1/7? I think I would use the identity 1-sin^2 theta to replace cos2theta, but I'm not sure.

zepdrix (zepdrix):

\(\Large \cos^2\theta \qquad\text{or}\qquad \cos(2\theta)\) ?

OpenStudy (anonymous):

Cos (2theta)

OpenStudy (john_es):

\[\cos(2\theta)=\cos^2\theta-\sin^2\theta=1-2\sin^2\theta\Rightarrow\] \[\Rightarrow \cos(2\theta)=1-2(1/7)^2=47/49\]

OpenStudy (anonymous):

Thank you